# PARCv2: Physics-aware Recurrent Convolutional Neural Networks for Spatiotemporal Dynamics Modeling

### v2

@@ A.2 Navier-Stokes equation
The flow around an obstacle has been a standard benchmark problem in fluid mechanics for decades, with laminar vortex shedding emerging as a particularly interesting and extensively studied phenomenon. Flow separation and vortex shedding occur due to the high positive pressure gradient, resulting in flow fields characterized by relatively sharp features that evolve rapidly in space, especially at high Reynolds numbers. In comparison to the Burgers’ equation case, the flow around an obstacle exhibits more advective characteristics, posing additional challenges for machine learning, as discussed earlier. Therefore, we selected this benchmark problem to assess and validate the modeling capability of PARCv2, building on its success with the Burgers’ equation.
@@
Oftentimes, the magnitude of the Reynolds number $R$ is a significant influence on the stability of fluid. When $R$ is small, the flow is inert and sluggish, rendering a rather stable flow pattern. On the other hand, when $R$ is large, the flow becomes unstable and starts to render complex patterns.
In many physics-aware deep learning works in literature, a frequently used benchmark is how well the model can predict unstable fluid behavior with large Reynolds numbers. Similar to these works, we consider a cylindrical cross-section of 0.25 m diameter embedded in a 2 m $\times$ 1 m rectangular domain, as illustrated in Figure 6. The circular obstacle was positioned $0.5(m)$ far from the left and $0.5(m)$from the top edge. We assume the initial velocity field to be uniformly distributed at the left side of the domain with initial velocity $u_{0}=1(m/s)$. Here the fluid density, $\rho$, is derived as: $\rho=4\cdot Re(kg/(m^{3}))$. Meanwhile, the kinematic viscosity, $\nu$, is derived from the dynamic viscosity $\mu$ as $\nu=\frac{\mu}{\rho}$.
For the train-test split, the Reynolds number was varied according to Table 4. The training data was obtained using finite volume methods with ANSYS Fluent. The initial data was provided in the triangular mesh format and was rasterized to fit with the use of CNN using interpolation. The rasterized domain has the size of $128\times 256$ $(pixels)$ with each pixel corresponding to a $0.0078125\times 0.0078125$ $(m)$ area.
@@ A.4 Measuring prediction accuracy and solution quality
Here, the subscript $ij$ indicates quantities specific to the grid location $(i,j)$. The superscript $hs$ indicates quantities specific to hotspots. $A_{ij}$ is the area of the grid cell occupied in the hotspot region, which is constant since a uniform grid was employed.
A.5 PhyCRNet for spatiotemporal dynamics modeling
Detail architecture for each problem
During our experiment, we kept the architecture of PhyCRNet intact with the one presented in the paper Ren et al. [2022]: 5 convolution encoder blocks with filter size of [4, 3, 4, 3, 4] and striding of [2, 1, 2, 1, 2], 1 Conv-LSTM block with filter size of 3 and striding of 1, and 1 pixel shuffle layer with an upscaling factor of 8 as 3 2x downsampling were performed during the encoding process. To eliminate the effect of number of parameters on prediction accuracy, we picked the number of channels to be [32, 32, 64, 64, 128, 128] for 2D Burger’s equations and [128, 128, 256, 256, 512, 512] for Navier-Stokes equation to maintain parameter parity with PARCv2, giving us $\sim$1 million parameters for 2D Burgers and $\sim$20 million parameters for N-S equations.
Training procedure
For PhyCRNet, we tested the performance of the model with only MSE data loss, with only PDE loss, and with a weighted combination of MSE and PDE loss on the 2D Burger’s Equation problem, and we only used a weighted combination of MSE and PDE loss on the N-S equation experiment. We used Adam optimizer with an initial learn rate of $10^{-4}$ and for every 100 epochs we halve the learn rate with a minimal learn rate of $10^{-6}$. We stopped training when the test loss stopped improving for 50 epochs.
A.6 FNO/PIFNO for spatiotemporal dynamics modeling
Detail architecture for each problems
In our experiments, we maintained the same architecture of Fourier Neural Operator (FNO) as the FNO-2d presented in the Li et al. [2020]: 2d FNO and a RNN structure in time. The 2d FNO first lifts the input to 32 channels for 2D Bugers (96 for N-S equations, 72 for EM) and then goes through 5 Fourier layers with 32 hidden channels (96 for N-S equations, 72 for EM), and lastly projects down to the number of channels in the output. The number of modes included in each Fourier layer is 15 for 2D Bugers and 24 for both N-S equations and EM, giving us $\sim$1 million parameters for 2D Burgers, $\sim$20 million parameters for N-S equations and $\sim$15 million parameters for EM, maintaining parameter parity with all other models presented in this work.
Training procedure

### v3

@@ A.2 Navier-Stokes Equation
Fluid flow around an obstacle has been a standard benchmark problem in fluid mechanics for decades, with laminar vortex shedding emerging as a particularly interesting and extensively studied phenomenon. Flow separation and vortex shedding occur due to high positive pressure gradients, resulting in flow fields characterized by relatively sharp features that evolve rapidly in space, especially at high Reynolds numbers. In comparison to the Burgers’ equation case, the flow around an obstacle exhibits more advective characteristics, posing additional challenges for machine learning, as discussed earlier. Therefore, we selected this benchmark problem to assess and validate the modeling capability of PARCv2, building on its success with the Burgers’ equation.
@@
Oftentimes, the magnitude of the Reynolds number $R$ has a significant influence on the stability of fluid. When $R$ is small, the flow is inert and sluggish, rendering a rather stable flow pattern. On the other hand, when $R$ is large, the flow becomes unstable and starts to render complex patterns.
In many physics-aware deep learning works in literature, a frequently used benchmark is how well the model can predict unstable fluid behavior with large Reynolds numbers. Similar to these works, we consider a cylindrical cross-section of 0.25 m diameter embedded in a 2 m $\times$ 1 m rectangular domain, as illustrated in Figure 6. The circular obstacle was positioned $0.5$ m far from the left and $0.5$ m from the top edge. We assume the initial velocity field to be uniformly distributed at the left side of the domain with initial velocity $u_{0}=1$ m/s. Here the fluid density, $\rho$, is derived as: $\rho=4\cdot R$ kg/m3. Meanwhile, the kinematic viscosity, $\nu$, is derived from the dynamic viscosity $\mu$ as $\nu=\frac{\mu}{\rho}$.
For the train-test split, the Reynolds number was varied according to Table 4. The training data was obtained using finite volume methods with ANSYS Fluent. The initial data was provided in the triangular mesh format and was rasterized to fit with the use of CNN using interpolation. The rasterized domain has the size of $128$ pixels × $\times\,256$ pixels with each pixel corresponding to a $7.8125\,\mu$m ×$\times 7.8125\,\mu$m area.
@@ A.4 Measuring Prediction Accuracy and Solution Quality
In Eqs. eqs. 27, 28, 29, 30, 31 and 32, the subscript $ij$ indicates quantities specific to the grid location $(i,j)$. The superscript $hs$ indicates quantities specific to hotspots. $A_{ij}$ is the area of the grid cell occupied in the hotspot region, which is constant since a uniform grid was employed.
A.5 PhyCRNet
During our experiment, we kept the architecture of PhyCRNet intact with the one presented in the paper Ren et al. [2022]. Five convolution encoder blocks with filter sizes [4, 3, 4, 3, 4] and strides [2, 1, 2, 1, 2], one Conv-LSTM block with filter size 3 and stride 1, and one pixel shuffle layer with an upscaling factor of 8 as three 2x downsampling were performed during the encoding process. To eliminate the effect of number of parameters on prediction accuracy, we picked the number of channels to be [32, 32, 64, 64, 128, 128] for 2D Burger’s equations and [128, 128, 256, 256, 512, 512] for Navier-Stokes equations to maintain parameter parity with PARCv2, yielding $\sim$1 million parameters for 2D Burgers and $\sim$20 million parameters for N-S equations.
For PhyCRNet, we tested the performance of the model with only MSE data loss, with only PDE loss, and with a weighted combination of MSE and PDE loss on the 2D Burger’s Equation problem, and we only used a weighted combination of MSE and PDE loss on the N-S equation experiment. We used Adam optimizer with an initial learning rate of $10^{-4}$ and for every 100 epochs we halved the learning rate with a minimal learning rate of $10^{-6}$. We stopped training when the test loss stopped improving for 50 epochs.
A.6 FNO/PIFNO
In our experiments, we maintained the same architecture of the Fourier Neural Operator (FNO) as the FNO-2d presented in the Li et al. [2020]: 2d FNO and an RNN structure in time. The 2d FNO first lifts the input to 32 channels for 2D Burgers (96 for N-S equations, 72 for EM) and then goes through 5 Fourier layers with 32 hidden channels (96 for N-S equations, 72 for EM), and lastly projects down to the number of channels in the output. The number of modes included in each Fourier layer is 15 for 2D Burgers and 24 for both N-S equations and EM, giving us $\sim$1 million parameters for 2D Burgers, $\sim$20 million parameters for N-S equations and $\sim$15 million parameters for EM, maintaining parameter parity with all other models presented in this work.
@@ A.6 FNO/PIFNO
A.7 Detailed Implementation of PARCv2
The detailed implementation of PARCv2 and necessary data for reproducing the results can be found in https://github.com/hphong1990/PARCv2.
For PARCv2, we used Adam optimizer with a learning rate of $10^{-5}$. As mentioned before, we trained PARCv2 using a two-step training. The differentiator was trained first for 500 epochs. Consequently, the differentiator is frozen and the integrator was trained for another 500 epochs. For the PARC setting without the data-driven integrator, only the differentiator was trained. No weight decay or any extra regularization was applied.
A.8 Comparison of Computational Cost
Table 5 reports the computation cost of PARCv2 and other baselines compared to DNS. All baselines were tested on an NVIDIA A100 Graphic Computing Unit (GPU). DNS was conducted on a high-performance computing (HPC) system with Intel® Xeon E5-2699v4 CPU @ 2.80 GHz (1,320 processors) and NVIDIA P100 GPUs accelerator.
Table 5: Computation cost comparison
Burgers Naiver-Stokes EM
PhyCRNet 0.264 (s) 0.713 (s) --- -
FNO 0.433 (s) 0.939 (s) 0.153 (s)
PIFNO 0.433 (s) 0.939 (s)
PARCv2 0.445 (s) 0.278 (s) 0.820(s)
PARC (num. int.) 0.122 (s) 0.127 (s) 0.089 (s)
PARC (NN. int.) 0.158 (s) 0.090 (s) 0.149 (s)
DNS 493.2 (s) 1-2 (hours) 2-3 (hours)